An aqueous solution of ammonium hydrogen phosphate has a concentration of 1.350 M and a density of 1.045 g/mL. (a) What is the percent by mass of the ammonium hydrogen phosphate in the solution?












(b) How many atoms of hydrogen would be in the ammonium hydrogen phosphate dissolved in 2.55 mL of the solution?
1
Expert's answer
2018-12-08T02:25:10-0500

Answer:

a) 1.350 M = 1.350 moles (NH4)2HPO4/L solution.

1.350 moles (NH4)2HPO4 = 1.350*(132.0562 g/mole) = 178.27 g

mass of 1000 ml of the solution is 1000*1.045 g/ml = 1045 g and 178.27 g is (NH4)2HPO4 therefore

w = mass of solute/mass of solution

w = (178.27/1045) * 100 = 17 %

b) 1.350 mole - 1000 ml

x mole - 2.55 ml

x = 0.003 mole (NH4)2HPO4

1 mole (NH4)2HPO4 - 9 mole of H

0.003 mole - x mole of H

x = 0.027 mole of H

N = Na*n, where Na is Avogadro constant, n is the amount of substance

N = 6.02*10^23 * 0.027 = 1.62*10^22